How do I get the coordinates of the vertex of the graph of Y=x^2+2x-3
Answer by scott8148(6628)

for y-k = a(x-h)^2 , the vertex is (h,k)

y+3 = x^2 + 2x

completing the square ___ y+3+1 = x^2 + 2x + 1 ___ y+4 = (x+1)^2

vertex is ___ (-1,-4)